c...
If the sides of a square are decreased by 3cm, the area is decreased by 81cm^2.What were the dimensions of the original square?

Answer by ankor@dixie-net.com(22740)

If the sides of a square are decreased by 3cm, the area is decreased by 81cm^2.What were the dimensions of the original square?
:
Let 1 side of the original square = x
Then the area of the original square = x^2
:
The side of the smaller square would = (x-2); original side decreased by 2cm
Area of the small square = (x-2)^2
:
Problem states:
Original square area - smaller square area = 81 sq/cm
x^2 - (x-2)^2 = 81
:
FOIL (x-2)(x-2)
x^2 -(x^2 - 4x + 4) = 81
:
Remove brackets, remember to change the signs inside the brackets:
x^2 - x^2 + 4x - 4 = 81
:
conveniently, the x^2's are eliminated:
4x = 81 + 4
4x = 85
x = 85/4
x = 21.25 cm is the side of the original square:
:
The smaller square would be 21.25 - 2 = 19.25
:
Check our solution by finding the area of each square:
21.25^2 = 451.5625
19.25^2 = 370.5625
-------------------
Differ by 81.00 sq/cm as given
:
Could you follow this OK? Any questions?